# Floating-Base Deep Lagrangian Networks

### v1

@@ VI-C Real Robots
To assess all methods also on real data, we ran tests with 4 different robots: HyQReal2, Spot, Spot+Arm, and Talos.
As a ground truth model is not available, we need to estimate $\boldsymbol{\tau}_{\nu}$, which in the case of a legged robot is
@@
where $\boldsymbol{\tau}_{\mathrm{q}}\in\mathbb{R}^{\mathrm{n}_{\mathrm{q}}}$ are the joint torques; $\mathbf{J}_{\mathrm{c}}{}_{i}$ and $\mathbf{f}_{\mathrm{c}}{}_{i}$ are the $i$th contact Jacobian and force. All evaluated robots are equipped with torque sensors. For contact forces, Talos uses an ankle sensor, HyQReal2 relies on a force estimator based on its dynamical model, and Spot’s forces were measured with external 6D force plates. For HyQReal2 and Talos, trajectories similar to Go2 in simulation were executed, producing 75k and 125k samples, respectively. To keep Spot on the force plates, small velocity commands were applied while trotting in place, resulting in 66k samples. For the arm, sine-wave trajectories with uniformly sampled parameters were tracked by a position controller, yielding 45k samples.
Although the estimation of $\boldsymbol{\tau}_{\nu}$ requires information about the robot’s kinematics, all the methods still learn agnostic to any kinematic parameter.
Table III reports the NMSE of the different methods across robots. Unlike in simulation, real robot dynamics are affected by actuator and contact dynamics, friction, noise, state estimation errors, and other unmodeled effects. Consequently, DNEA methods struggle to accurately estimate the inverse dynamics torques due to limited model complexity, being outperformed by the FFNN. In contrast, DeLaN-PP, FeLaN-BS, and FeLaN achieved the lowest $r\text{NMSE}$, highlighting the effectiveness of the proposed potential energy parametrization. The results hightlight the advantage of FeLaN, which achieved similar performance with a more interpretable and physical consistent model, preserving properties such as stationary mass and branch-induced sparsity.
TABLE III: Inverse dynamics normalized mean squared error for the real robot data.
Method HyQReal2 Spot Spot with Arm Talos $r\text{NMSE}$
Train Test Train Test Train Test Train Test
FFNN $0.2\pm 16.1 11$$1.5\pm 18.0$ $\mathbf{5.4\pm 0.2}$ $\mathbf{4.0\pm 0.3}$ $\mathbf{7.8\pm 0.6}$ $\mathbf{9.5\pm 0.3}$ $28.0\pm 0.1$ $27.0\pm 0.1$ $0.210$
DeLaN $88.3\pm 5.8$ $278.0\pm 167.0$ $6.5\pm 0.2$ $\mathbf{4.0\pm 0.2}$ $20.2\pm 12.0$ $10.9\pm 0.9$ $\mathbf{4.2\pm 0.1}$ $\mathbf{4.5\pm 0.4}$ $0.268$
DeLaN-$P$ $\mathbf{8.7\pm 0.9}$ $\mathbf{10.1\pm 1.1}$ $7.4\pm 0.1$ $4.9\pm 0.3$ $3.1\pm 0.2 11$$1.1\pm 0.2$ $9.0\pm 0.5$ $8.6\pm 0.4$ $\mathbf{0.102}$
FeLaN-BS $10.7\pm 0.8$ $12.6\pm 0.7$ $7.8\pm 0.1$ $4.9\pm 0.2$ $12.4\pm 0.3$ $10.9\pm 1.1$ $9.5\pm 0.2$ $8.9\pm 0.2$ $0.106$
FeLaN $10.5\pm 0.2$ $14.8\pm 0.8$ $8.4\pm 0.9$ $5.1\pm 0.3$ $4.0\pm 0.5 11$$1.4\pm 0.6$ $9.6\pm 0.4$ $9.1\pm 0.4$ $0.126$
DNEA NS $16.0\pm 0.3$ $20.3\pm 0.3$ $11.1\pm 0.1$ $6.0\pm 0.1$ $17.6\pm 0.2$ $13.1\pm 0.2$ $18.5\pm 0.8$ $16.9\pm 0.7$ $0.271$
DNEA PD $16.2\pm 0.2$ $20.4\pm 0.4$ $11.2\pm 0.1$ $6.0\pm 0.1$ $17.7\pm 0.1$ $13.2\pm 0.2$ $19.6\pm 1.2$ $18.0\pm 1.0$ $0.287$
DNEA Tri $16.2\pm 0.2$ $20.7\pm 0.4$ $11.2\pm 0.1$ $6.0\pm 0.1$ $17.8\pm 0.1$ $13.3\pm 0.2$ $19.9\pm 0.7$ $18.2\pm 0.6$ $0.290$
DNEA Cov $16.3\pm 0.2$ $20.7\pm 0.3$ $11.3\pm 0.1$ $6.1\pm 0.1$ $17.8\pm 0.1$ $13.3\pm 0.2$ $20.0\pm 1.1$ $18.4\pm 1.0$ $0.294$
DNEA Log $16.4\pm 0.2$ $20.8\pm 0.3$ $11.3\pm 0.1$ $6.0\pm 0.1$ $17.8\pm 0.1$ $13.3\pm 0.2$ $20.0\pm 0.9$ $18.4\pm 0.7$ $0.293$
@@ VII Conclusion
In this work, we introduced a novel physically consistent parametrization of the floating-base inertia matrix, based on a reordered Cholesky factorization, which achieves branch-induced sparsity and ensures full physical consistency of the composite spatial inertia. Building on this parametrization, we proposed FeLaN, a grey-box method for physically consistent system identification of floating-base dynamics. FeLaN learns the whole-body dynamics model using only prior knowledge on the kinematic chains, without requiring any knowledge of dynamics or other kinematic parameters. We validated FeLaN on both simulated and real robot data, showing performance comparable to state-of-the-art methods while offering greater interpretability.
@@
APPENDIX
VII-A Training
The datasets are split into 90% for training and 10% for testing. Table IV lists the used hyperparameters.
TABLE IV: Training hyperparameters.
Parameter Value Parameter Value
Activation Tanh $\epsilon_{\mathrm{L}}$ 0.01
Batch Size 1024 $\epsilon_{\mathrm{m}}$ 0.1
Learning Rate $5\times 10^{-4}$ $\epsilon_{\mathrm{D}}$ 0.01
Weight decay $10^{-5}$ Gradient Clip Norm 1000

### v2

@@ VI-C Real Robots
To evaluate all methods on real-world data, we ran experiments on four robots: HyQReal2, Spot, Spot with Arm, and Talos. As a ground truth model is not available, we estimate $\boldsymbol{\tau}_{\nu}$, which for a legged robot with $\mathrm{n}_{\mathrm{c}}$} contacts is given by
@@
where $\boldsymbol{\tau}_{\mathrm{q}}\in\mathbb{R}^{\mathrm{n}_{\mathrm{q}}}$ are the joint torques; $\mathbf{J}_{\mathrm{c}}{}_{i}$ and $\mathbf{f}_{\mathrm{c}}{}_{i}$ are the $i$th contact Jacobian and force. All evaluated robots have torque sensors. For contact forces, Talos uses an ankle sensor, HyQReal2 relies on a force estimator based on its dynamical model, and Spot’s forces were measured with external 6D force plates. Although the estimation of $\boldsymbol{\tau}_{\nu}$ requires kinematics information, all the methods still learn agnostically to any kinematic parameter. For HyQReal2 and Talos, trajectories similar to those in simulation were executed, producing 75k and 125k samples, respectively. To keep Spot on the force plates, small velocity commands were sent while trotting in place, resulting in 60k samples. For the arm, a position controller tracked sine-wave trajectories with uniformly sampled parameters, yielding 45k samples.
Table III reports the NMSE results, including the weighted average $\overline{r\text{NMSE}}$ over real and simulated robots. Unlike in simulation, real robot dynamics are affected by actuator and contact dynamics, noise, and other unmodeled effects. Consequently, DNEA methods struggle to accurately estimate the torques due to limited model complexity, being in general outperformed by most of the other methods. Following the simulation results on Talos, the performance gap of DNEAs is larger for the heavier robots. On HyQReal2 and Talos, 125 kg and 100 kg, respectively, DNEA errors are roughly three times higher than the best. Their larger mass implies larger contact forces and thereby larger torques than on Spot. This tends to amplify the influence of actuator and contact dynamics that the models do not capture, thereby limiting prediction accuracy. In contrast, DeLaN and MLP achieve the best robot-specific results. Because these methods impose weaker or no physical constraints, they are more flexible in learning input–output mappings that may not be physically consistent. However, they demonstrated sensitivity to initialization and hyperparameter choices, and do not perform consistently across all robots.
Overall, DeLaN-PP, FeLaN-BS, and FeLaN remain competitive with the best robot-specific models while achieving the best $r\text{NMSE}$. This supports incorporating physics into deep learning by improving consistency across robots, reducing sensitivity to dataset-specific biases. In particular, it highlights the effectiveness of our potential energy parameterization. Considering $\overline{r\text{NMSE}}$, FeLaN achieves the best overall performance across simulation and real robots, followed by FeLaN-BS, underscoring the benefit of branch-induced sparsity and full physical consistency of the composite spatial inertia.
@@ VII Conclusion
In this work, we introduced a novel physically consistent parametrization of the floating-base inertia matrix, based on a reordered Cholesky factorization, which achieves branch-induced sparsity and ensures full physical consistency of the composite spatial inertia. Building on this parametrization, we proposed FeLaN, a grey-box method for physically consistent system identification of floating-base dynamics. FeLaN learns the whole-body dynamics model using only prior knowledge on the kinematic chains, without requiring any knowledge of dynamics or other kinematic parameters. We validated FeLaN on both simulated and real robot data, showing overall better performance than state-of-the-art methods while offering greater interpretability.